Elimination
8s +12t=16
6s + 9t=12

Answer by Boreal(15235)

8s +12t=16
6s + 9t=12; multiply by -4/3
-8s-12t=-16
add them and 0=0
These are the same line. Infinite solutions, not independent. All are appropriate terms.